# Rational angle bisectors on the coordinate plane and solutions of Pell’s equations

### v18

@@ 1 Introduction
For which rational numbers are the slopes of the angle bisectors between two straight lines with slopes rational?
@@
Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without errors due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
@@
Let be an integral solution of We say that is strictly primitive if are coprime.
Let be positive integral solutions of We say is smaller than if or “ and ”
For each equation we call its minimum positive integral solution its fundamental solution.
@@
Suppose that is a square number, and let be an integral solution of If is primitive, that is, are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that are coprime and is square-free. In particular, every integral solution of is strictly primitive.
@@
Furthermore, for each where we denote the conjugate, norm of by
@@
Suppose that are the -components of rational solutions of the equation Then is given by
(1.3) for some such that
Suppose that are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
(1.4) for some such that and In addition, can be written in the form
with some satisfying
@@
(1.5) for some for each prime number where “ and ” if In this case, its strictly primitive integral solution satisfies
(1.6) for some such that
(1.7) and
@@
(1.8) for some where (1.7) and
(1.9) hold.
@@ 2 Rational solutions of Pell’s equations
Conversely, if has a rational solution with then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
@@
Let If have integral solutions respectively, then has an integral solution
@@
Let If has an integral solution, then there exists a finite number of its integral solutions such that its every integral solution satisfies
@@
For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then for some
@@
and therefore Hence
Thus and therefore the solution is not strictly primitive.
Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then
@@
Hence
Thus and therefore the solution is not strictly primitive.∎
@@
We prove the contraposition. Suppose that has no strictly primitive integral solutions. Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that
@@
Case 2: Suppose that Then In this implies and therefore since Hence
Thus This implies that has no strictly primitive integral solutions. ∎
@@
Let be the monoid of all ideals in For each we define if
@@
by the finiteness of where are coprime since is primitive. Furthermore, are coprime, since and These imply that has a strictly primitive integral solution.
@@
In these cases, is not strictly primitive. Thus the desired assertion holds. ∎
@@
Unless “ and ” then is not greater than the ideal class number of for each
@@
Suppose that Then the pair of odd integers defined by satisfies This implies that are coprime, and therefore and In this case, holds by the minimality of
@@
If or “ and ” or then consists of all prime numbers which split in
@@
Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
@@
in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore which contradicts that are coprime. For each we can see that is divisible by either or by the same argument as above, which implies
@@
Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1), (4). The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that are coprime. The ideal is decomposed as
with in The ideal
@@
This implies and therefore which contradicts that are coprime.
@@
(2.7) for some satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
@@
The equation has the fundamental solution however, has no integral solutions. The fundamental unit of is The equations have the fundamental solutions respectively. Every integral solution of satisfies one of the following conditions for some
@@
for some such that
@@
Let be a rational solution of where are coprime integers and Then is an integral solution of and satisfies
for some where
@@ 3 Rational angle bisectors
with Substituting these into ( ‣ 1) and multiplying both sides by we obtain
For any prime number the parities of and coincide with each other, since are square numbers. Let be the product of every prime number such that these valuations are odd (define if there are no such prime numbers). Then there exist such that
where are the -components of the integral solutions of and are the -components of the rational solutions of with Substituting (3.2) into (3.1) and dividing both sides by we obtain
or equivalently
@@
Suppose that are the -components of rational solutions of respectively. Let
with Then we have
@@
Suppose that are the -components of rational solutions of for some square-free integer respectively. In general, if and then
@@
In Theorem 2, for such that has a rational solution and the values of are summarized in the following table. For reference, we add the values of the ideal class number the fundamental unit of and below the values of respectively.

### v20

@@ 1 Introduction
For which rational numbers and are the slopes of the angle bisectors between two straight lines with slopes and rational?
@@
Essentially, Problem 1 has the meaning when the bisectors of can be drawn by connecting and other lattice points for given lattice points and which is important in drawing techniques. Furthermore, in engineering, we can specify the radiation range and the axis of light with a ratio of integers without errors due to approximations to irrational numbers by using a solution to this problem. Problem 1 is reduced to solving the equation
@@
Let be an integral solution of We say that is strictly primitive, if and are coprime.
Let be positive integral solutions of We say is smaller than if or if and
For each equation of and we call its minimum positive integral solution its fundamental solution.
@@
Suppose that is a square number, and let be an integral solution of If is primitive, that is, and are coprime, then is strictly primitive; otherwise, there exists a common prime divisor of and dividing twice and also twice, which contradicts that and are coprime, and is square-free. In particular, every integral solution of is strictly primitive.
@@
Furthermore, for each where we denote the conjugate and norm of by
@@
Suppose that and are the -components of rational solutions of the equation Then is given by
(1.3) for some such that and
Suppose that and are the -components of rational solutions of a common negative Pell’s equation for some square-free integer Then is given by
(1.4) for some such that and In addition, and can be written in the form
for some and satisfying
@@
(1.5) for some for each prime number where and if In this case, its strictly primitive integral solution satisfies
(1.6) for some and such that
(1.7) -
@@
(1.8) for some and where
(1.9) and (1.7) hold.
@@ 2 Rational Solutions of Pell’s Equations
Conversely, if has a rational solution with and then which implies that has no prime divisors congruent to modulo by Fermat’s theorem as above, and so does ∎
@@
Let If and have integral solutions and respectively, then has an integral solution
@@
Let Suppose that has an integral solution. Then, for this equation, there exists a finite number of integral solutions such that every integral solution satisfies
@@
For each let be an integral solution of and let Suppose that the integral solution of defined by (2.1) is not strictly primitive. Then and for some
@@
and therefore Hence and
Thus, in each case, and which imply that the solution is not strictly primitive.
Let be an integral solution of Suppose that the solution of defined by (2.2) is not strictly primitive. Then and
@@
Hence and
Thus, in each case, and which imply that the solution is not strictly primitive.∎
@@
We prove the contraposition. Suppose that has no strictly primitive integral solutions. Let be an integral solution of Then defined by is an integral solution of which implies that there exists a prime divisor of such that and
@@
Case 2: Suppose that Then In this implies and therefore since Hence and
Thus, in each case, and which imply that has no strictly primitive integral solutions. ∎
@@
Let be the monoid of all nonzero ideals in For each we define if
@@
by the finiteness of where and are coprime since is primitive. Furthermore, and are coprime, since and These imply that has a strictly primitive integral solution.
@@
Thus, in each case, is not strictly primitive. This proves the desired assertion. ∎
@@
Unless and then is not greater than the ideal class number of for each
@@
Suppose that Then the pair of odd integers defined by satisfies This implies that and are coprime, and therefore and In this case, holds by the minimality of
@@
If or and or then consists of all prime numbers which split in
@@
Let and let denote the integer ring of For each prime number we take a prime ideal in lying over in and let Note that and are primitive for each if splits in Let be a strictly primitive integral solution of Let be the set of all prime divisors of and let For each let and be the quotient and remainder, respectively, when dividing by For each let The equality can be expressed as with ideals in which implies
@@
in by Theorem 7, and therefore the ideal is divisible by either or ; otherwise, in and also by (2.4), and therefore and which contradict that and are coprime. For each we can see that is divisible by either or by the same argument as above, which implies
@@
Assume that Then is unramified in by Theorem 8, and by Lemma 1 (1) and (4). The prime ideal divides both and only once by its self-conjugacy; otherwise, in and therefore in by (2.6), which contradicts that and are coprime. The ideal is decomposed as
with and in The ideal
@@
This implies and therefore and which contradict that and are coprime.
@@
Subcase 2-2: Suppose that Then is unramified in by Lemma 1 (3), and and by Lemma 1 (2) and (4). The ideal is decomposed as
with and in The ideal
Letting for each and we obtain (1.6).
@@
(2.7) for some and satisfying (1.7) and the condition obtained by replacing with in (1.5) because of (1). If then (1.9) holds, since If then Multiplying both sides of (2.7) by
@@
The equation has the fundamental solution however, has no integral solutions. The fundamental unit of is The equations and have the fundamental solutions and respectively. Every integral solution of satisfies one of the following conditions for some
@@
for some and such that
@@
Let be a rational solution of where and are coprime integers and Then is an integral solution of and satisfies
for some and where
@@ 3 Rational Angle Bisectors
with and Substituting these into ( ‣ 1) and multiplying both sides by we obtain
For any prime number the parities of and coincide with each other, since and are square numbers. Let be the product of every prime number such that these valuations are odd (define if there are no such prime numbers). Then there exist such that
where and are the -components of the integral solutions of Furthermore, and are the -components of the rational solutions of with and Substituting (3.2) into (3.1) and dividing both sides by we obtain
or equivalently,
@@
Suppose that and are the -components of rational solutions and of respectively. Let
with and Then we have
@@
which implies
Suppose that and are the -components of rational solutions and of for some square-free integer respectively. In general, if and with then
which imply
@@
In Theorem 2, for such that has a rational solution and the values of and are summarized in the following table. For reference, we add the values of and below the values of and respectively. We also add the value of the ideal class number of